# Universal quantum control over non-Hermitian continuous-variable systems

### v1

@@ III.1 Model and Hamiltonian
Figure 1: Sketch of an open cavity magnonic system under control. More than the coherent exchange coupling with the strength $J$ and the phase $\varphi$ between the cavity mode $a$ and the magnon mode $b$, the cooperative coupling of the two-mode system to the environment gives rise to the gain or loss rate $\gamma_{a}$ of the cavity mode, the loss rate $\gamma_{b}$ of the magnon mode, and the dissipation coupling $i\Gamma e^{i\Theta}$ between them. $\Theta$ can be 0 or $\pi$ in experiments Wang et al. (2019a).
In this section, we use our UQC theory for the general non-Hermitian bosonic systems in Sec. II to analyze an open two-mode bosonic system that consists of a single-mode cavity and a yttrium iron garnet sphere in the Kittel mode Harder et al. (2018); Grigoryan et al. (2018); Bhoi et al. (2019); Yang et al. (2019); Yao et al. (2019); Yu et al. (2019); Boventer et al. (2020) as shown in Fig. 1. The cavity and magnon modes are represented by the annihilation operators $a$ and $b$, respectively. They are mutually interacted by the time-dependent exchange coherent coupling $J(t)$ with the phase $\varphi$. The cooperative coupling between the two-mode system and the environment Metelmann and Clerk (2015); Wang et al. (2019a) leads to the gain or loss rates $\gamma_{a}$ and $\gamma_{b}$ of the cavity mode $a$ and the magnon mode $b$, respectively, as well as a dissipative coupling $i\Gamma e^{i\Theta}$ between them. Then the whole system can be described by a non-Hermitian Hamiltonian Metelmann and Clerk (2015); Wang et al. (2019a):
$H(t)=(\omega_{a}-i\gamma_{a}e^{i\varphi_{a}})a^{\dagger}a+[\omega_{b}(t)-i\gamma_{b}]b^{\dagger}b$ (20)
@@
Here $\omega_{a}$ is the static frequency of the cavity mode. $\omega_{b}(t)$ is the time-dependent frequency of the magnon modes, which can be flexibly tuned by the external bias magnetic field $B(t)$ Zhang et al. (2014); Zare Rameshti et al. (2022); Jin and Jing (2024). $\gamma_{b}>0$ is the fixed loss rate of magnon mode. $\gamma_{a}>0$ represents the gain or loss effect on the cavity mode when $\varphi_{a}=\pi$ or $\varphi_{a}=0$, respectively Zhang et al. (2017, 2025). The relative phase $\Theta=0$ or $\Theta=\pi$ between the dissipative coupling and the coherent coupling is determined by the propagation direction of the traveling waves Wang et al. (2019a). For simplicity, we set $\Theta=0$ in the following. The details about the derivation of the non-Hermitian Hamiltonian (20) with $\varphi_{a}=0$ through the Lindblad master equation can be found in appendix B, despite both $\varphi_{a}$ and $\Gamma$ can have many choices Wang et al. (2019a); Zhang et al. (2017, 2025).
@@
become real-valued, i.e., $E_{\pm}=\pm\sqrt{J^{2}-\gamma_{a}^{2}}$, when $H(t),\mathcal{PT}]=0$ and $J>\gamma_{a}$. Otherwise, they remain complex and the Hamiltonian is in the $\mathcal{PT}$-symmetry-broken phase. Whether holding $\mathcal{PT}$-symmetry or not, the system can approach EP under the conditions of $\Delta(t)=0$, $\varphi=\pi/2$, $\gamma_{a}=\gamma_{b}$, and $J=\pm\Gamma$ (when $\varphi_{a}=0$) or $J=\pm\sqrt{\Gamma^{2}+\gamma_{a}^{2}}$ (when $\varphi_{a}=\pi$), where the eigenenergies coalesce as $E_{+}=E_{-}$.
Conventionally, the control protocols proposed for non-Hermitian systems are based on their spectrum characteristics Ashida et al. (2020); Bender and Hook (2024), focusing on $\mathcal{PT}$-symmetry and EPs. In contrast, our UQC theory in Sec. II is essentially proposed in the instantaneous frames, which are irrespective to the spectral properties. In the following section, we will demonstrate that an arbitrary state can be perfectly transferred from the cavity mode to magnon mode, regardless of whether the system in the unbroken or broken phases of $\mathcal{PT}$-symmetry and whether EPs are present or not.
@@ III.2 Universal Heisenberg passages of two operators
With Eq. (21), we construct the non-unitary universal passages for the non-Hermitian cavity-magnonic system, where the final state can be automatically normalized. According to Eqs. (42) and (43), the time-dependent ancillary operators of an arbitrary two-mode system can be written as
@@
where the parameters $\theta(t)$ and $\alpha(t)$ manipulate the population and the relative phase of both cavity mode $a$ and magnon mode $b$, respectively. From Eqs. (45) and (46), the unitary transformation $\mathcal{V}(t)$ in Eq. (6) which acts as $\mathcal{V}^{\dagger}(t)\mu_{1}(t)\mathcal{V}(t)\rightarrow\mu_{1}(0)$ and $\mathcal{V}^{\dagger}(t)\mu_{2}(t)\mathcal{V}(t)\rightarrow\mu_{2}(0)$, takes the form of
@@
It is interesting to found that the preceding transformation from the time-dependent ancillary operators to the time-independent ones can also be formally implemented by $V_{\theta}(t)V_{\alpha}(t)]^{\dagger}\mu_{k}(t)V_{\theta}(t)V_{\alpha}(t)\rightarrow\mu_{k}(0)$, $k=1,2$, despite that $V_{\alpha}(t)$ and $V_{\theta}(t)$ are noncommutative, i.e., $V_{\alpha}(t),V_{\theta}(t)]\neq 0$.
@@
In parallel, for the system dynamics governed by the Hermitian conjugate $H^{\dagger}(t)$ in the dual space, the same conditions in Eq. (27) can activate the ancillary operator $\mu_{2}^{\dagger}(t)$ as the nonadiabatic passage. In particular, the time evolution takes the form of Eq. (19), where the global phase $\dot{f}_{2}(t)=-i[\gamma_{a}\exp(i\varphi_{a})+\gamma_{b}]-\dot{f}_{1}(t)$ with $\dot{f}_{1}(t)$ given by Eqs. (29) and (30). Similar to the passage $\mu_{1}^{\dagger}(t)$, a flexible and perfect state transfer can be completed along the passage $\mu_{2}^{\dagger}(t)$ under appropriate choices of $\theta(t)$ and $\alpha(t)$. The probability conservation can also be ensured in the end of the evolution.
@@ IV.1 Perfect state transfer
We first consider the Hamiltonian in the phase of $\mathcal{PT}$-symmetry. In this case, $\varphi_{a}=\pi$, $\gamma_{a}=\gamma_{b}$, $\Gamma=0$, and $\Delta(t)=0$. Then the constraint on $J(t)$ in Eq. (27) becomes $J(t)=\dot{\theta}(t)/\sin[\varphi+\alpha(t)]$. Suppose the cavity mode and magnon mode initially prepared in the Fock state $|n=5\rangle$ and the vacuum state $|0\rangle$, respectively, i.e., $|\psi(0)\rangle=|5\rangle_{a}|0\rangle_{b}$. Using Eqs. (23), (24), and (28), the system passage can evolve as $\mu_{1}^{\dagger}(0)=a^{\dagger}\rightarrow e^{if_{r}(\tau)}\mu_{1}^{\dagger}(\tau)=e^{if_{r}(\tau)}b^{\dagger}$ under the setting of $\alpha(t)=0$ and
@@
Then the initial state of the two-mode system $|\psi(0)\rangle$
$|\psi(0)\rangle=|5\rangle_{a}|0\rangle_{b}=\frac{(a^{\dagger})^{5}}{\sqrt{5!}}|0\rangle_{a}|0\rangle_{b}=\frac{[\mu_{1}^{\dagger}(0)]^{5}}{\sqrt{5!}}|0\rangle_{a}|0\rangle_{b}$ (32)
is transferred to be
$\frac{[\mu_{1}^{\dagger}(\tau)]^{5}}{\sqrt{5!}}|0\rangle_{a}|0\rangle_{b}=\frac{(b^{\dagger})^{5}}{\sqrt{5!}}|0\rangle_{a}|0\rangle_{b}=|0\rangle_{a}|5\rangle_{b}=|\psi(\tau)\rangle,$ (33)
demonstrating a perfect Fock-state transfer from the cavity mode to the magnon mode.
Under the non-unique condition in Eq. (31), we find that the probability conservation of the system state can be restored at the end of the evolution in the presence or in the absence of exceptional points by appropriately choosing the gain or loss rates and the dissipative coupling strength. Particularly for the Hamiltonian of $\mathcal{PT}$-symmetry, the imaginary part of global phase in Eq. (30) is reduced to be
@@
To ensure the vanishing of the time integral over Eq. (34), one can choose the gain or loss rate and dissipative coupling strength as
@@
given $\theta(t)$ is set as a linear function of time as in Eq. (31). Here the coefficient $\lambda$ scales the magnitudes of these rates or strength.
@@
Figure 2 demonstrates perfect nonadiabatic passages $\mu_{1}^{\dagger}(t)$, i.e., $\mathcal{F}_{5,0}(0)=1$ and $\mathcal{F}_{0,5}(\tau)=1$, under a $\mathcal{PT}$-symmetric Hamiltonian. It is found that for both evolutions of (a) avoiding and (b) crossing EPs, the Fock states with conserved excitations other than the initial and target states are scarcely populated. For example, in Fig. 2(a), we have at most $\mathcal{F}_{4,1}(0.19\tau)=0.073$, $\mathcal{F}_{3,2}(0.40\tau)=0.031$, $\mathcal{F}_{2,3}(0.62\tau)=0.032$, and $\mathcal{F}_{1,4}(0.82\tau)=0.074$. The total fidelity $\sum_{n}\mathcal{F}_{n}\equiv\sum_{n_{1},n_{2}}\mathcal{F}_{n_{1},n_{2}}$, being equivalent to the trace of the two-mode system ${\rm Tr}[|\psi(t)\rangle\langle\psi(t)|]$, can represent the probability conservation or non-conservation under a non-Hermitian Hamiltonian. In both Figs. 2(a) and (b), it is found that $\sum_{n}\mathcal{F}_{n}(0<t<\tau)<1$ due to the gain effect of the cavity mode and the loss effect of the magnon mode. Nevertheless, $\sum_{n}\mathcal{F}_{n}(t=\tau)=1$ at the end of the passage. In Figs. 2(c) and (d), we present the dynamics of real and imaginary parts of eigenenergies $E_{\pm}$ due to Eq. (22), relevant to the avoiding and crossing EPs in the parametric setting, respectively. The eigenenergies in Fig. 2(c) verifies that the system dynamics avoids the coalescence of eigenenergies and states. Figure 2(d) indicates that the system crosses EPs when $t=0.18\tau$ and $t=0.82\tau$. It is hardly to find a clear relevance between the perfect state transfer and the existence of the exceptional points of the system spectrum.
@@
Similar to Figs. 2(a) and (b), Figs. 3(a) and (b) demonstrate the fidelity dynamics $\mathcal{F}_{n_{1},n_{2}}$ of various five-excitation states and their summation $\sum_{n}\mathcal{F}_{n}=\sum_{n_{1},n_{2}}\mathcal{F}_{n_{1},n_{2}}$ within the broken phase of $\mathcal{PT}$-symmetry in the absence and presence of EPs, respectively. Here we set $\varphi_{a}=0$ and $\gamma_{a}=\gamma_{b}$. Then the imaginary part of global phase in Eq. (30) becomes
@@
One can check that the gain or loss rate and dissipative coupling strength in Eq. (35) are still applicable to neutralize the imaginary phase under the parametric setting in Eq. (31) for $\theta$.
In both Figs. 3(a) and (b), it is found that the initial population on $|5\rangle_{a}|0\rangle_{0}$ can be completely transferred to $|0\rangle_{a}|5\rangle_{b}$, even when the other five-excitation states could be temporally yet significantly populated during the passage. The Fock states with more number of excitations in the target (magnon) mode are sequently populated until $\mathcal{F}_{0,5}(\tau)=1$ at the end of the time evolution. Note our wavefunction $|\psi(t)\rangle$ is nonunitary and never renormalized. For $0<t<\tau$, the population summation of individual states is not subject to the probability conservation, i.e., $\sum_{n}\mathcal{F}_{n}<1$ or $\sum_{n}\mathcal{F}_{n}>1$, due to the loss effects of both the cavity and magnon modes and the dissipative coupling between them. The peak value of the population summation is found to be $\sum_{n}\mathcal{F}_{n}=1.32$ when $t=0.72\tau$ in Fig. 3(a), around which the dissipative coupling between the cavity and magnon modes dominates the loss effects of both modes. In Fig. 3(a) in the absence of EPs, we have $\mathcal{F}_{4,1}(0.22\tau)=0.35$, $\mathcal{F}_{3,2}(0.34\tau)=0.34$, $\mathcal{F}_{2,3}(0.45\tau)=0.40$, $\mathcal{F}_{1,4}(0.61\tau)=0.55$, and $\mathcal{F}_{0,5}(0.90\tau)=1.07$. And in Fig. 3(b) in the presence of EPs, we have $\mathcal{F}_{4,1}(0.35\tau)=0.24$, $\mathcal{F}_{3,2}(0.50\tau)=0.32$, $\mathcal{F}_{2,3}(0.61\tau)=0.46$, $\mathcal{F}_{1,4}(0.72\tau)=0.74$ and $\mathcal{F}_{0,5}(0.89\tau)=1.40$. Under the conditions in Eq. (35), the state norm becomes unit at the end of the passage. Again the associated evolutions of the real and imaginary parts of the eigenenergies are presented in Figs. 3(c) and (d), respectively. As shown in Fig. 3(c), the eigenenergies do not coalesce during the whole passage, confirming the avoidance of EPs. In Fig. 3(d), an EP occurs at $t=0.27\tau$, also being irrelevant to the system dynamics.
@@
Far beyond the Fock state, the universal nonadiabatic passage enables perfect transfer of diverse states, including the binomial code state Michael et al. (2016), the coherent state, the cat state, and the thermal state. In Fig. 4(a), the cavity and magnon modes are initially prepared in the binomial code state $|\phi_{b}\rangle=(\sqrt{3}|2\rangle+|6\rangle)/2$ Michael et al. (2016) and the vacuum state $|0\rangle$, respectively. The performance of our protocol can be evaluated by $\mathcal{F}_{\phi_{b},0}=|\langle\phi_{b}|\langle 0|\psi(t)\rangle|^{2}$ and $\mathcal{F}_{0,\phi_{b}}=|\langle 0|\langle\phi_{b}|\psi(t)\rangle|^{2}$. It is found that the binomial code state in the cavity mode can be perfectly transferred to the magnon mode with a unit fidelity. Along the passage, the probability summation during $0<t<\tau$ is not conserved due to the cavity gain and dissipative coupling between the cavity mode and the magnon mode. For example, we have ${\rm Tr}(\rho)=0.631$ when $t=0.234\tau$ and ${\rm Tr}(\rho)=1.63$ when $t=0.782$. However, the probability conservation is restored as ${\rm Tr}(\rho)=1$ when $t=\tau$. Similarly, in Figs. 4(b), (c), and (d) for the cavity mode initially prepared as the coherent state $|\alpha\rangle$ with $\alpha=5$, the cat state $|{\rm cat}\rangle=(|\alpha\rangle+|-\alpha\rangle)/\sqrt{2}$ with $\alpha=5$, and the thermal state $\rho_{\rm th}$ with $\bar{n}=5$, respectively, it is found that all of them can be perfectly transferred to the target mode. Note for the last mixed state, $\mathcal{F}_{{\rm thermal},0}$ and $\mathcal{F}_{0,{\rm thermal}}$ are not complementary to each other in any situation.
In parallel, a quantum state initially prepared in the magnon mode can be perfectly transferred to the cavity mode by activating the passage $\mu_{2}^{\dagger}(t)$ in dual space. It can be simply realized under almost the same conditions as for the passage $\mu_{1}^{\dagger}(t)$, i.e., the parameters employed in Figs. 2, 3, and 4, by replacing $\varphi=\pi/2$ with $\varphi=-\pi/2$.

### v2

@@ III.1 Model and Hamiltonian
Figure 1: Sketch of an open cavity magnonic system under control. More than the coherent exchange coupling with strength $J$ and phase $\varphi$ between the cavity mode $a$ and the magnon mode $b$, the cooperative coupling of the two-mode system to the environment gives rise to the gain or loss rate $\gamma_{a}$ of the cavity mode, the loss rate $\gamma_{b}$ of the magnon mode, and the dissipation coupling $i\Gamma e^{i\Theta}$ between them. $\Theta$ can be 0 or $\pi$ in experiments Wang et al. (2019a).
In this section, we use our UQC theory for the general non-Hermitian bosonic systems in Sec. II to analyze an open two-mode bosonic system that consists of a single-mode cavity and a yttrium iron garnet sphere in the Kittel mode Harder et al. (2018); Grigoryan et al. (2018); Bhoi et al. (2019); Yang et al. (2019); Yao et al. (2019); Yu et al. (2019); Boventer et al. (2020) as shown in Fig. 1. The cavity and magnon modes are represented by the annihilation operators $a$ and $b$, respectively. They are mutually interacted by the time-dependent exchange coupling $J(t)$ with phase $\varphi$. The cooperative coupling between the two-mode system and the environment Metelmann and Clerk (2015); Wang et al. (2019a) leads to the gain or loss rates $\gamma_{a}$ and $\gamma_{b}$ of the cavity mode $a$ and the magnon mode $b$, respectively, as well as a dissipative coupling $i\Gamma e^{i\Theta}$ between them. Then the whole system can be described by a non-Hermitian Hamiltonian Metelmann and Clerk (2015); Wang et al. (2019a):
$H(t)=\left(\omega_{a}-i\gamma_{a}e^{i\varphi_{a}}\right)a^{\dagger}a+\left[\omega_{b}(t)-i\gamma_{b}\right]b^{\dagger}b$ (20)
@@
Here $\omega_{a}$ is the free frequency of the cavity mode. $\omega_{b}(t)$ is the time-dependent frequency of the magnon modes, which can be flexibly tuned by the external bias magnetic field $B(t)$ Zhang et al. (2014); Zare Rameshti et al. (2022); Jin and Jing (2024). $\gamma_{b}>0$ is the fixed loss rate of magnon mode. $\gamma_{a}>0$ represents the gain or loss effect on the cavity mode when $\varphi_{a}=\pi$ or $\varphi_{a}=0$, respectively Zhang et al. (2017, 2025). The phase $\Theta=0$ or $\Theta=\pi$ of the dissipative coupling is determined by the propagation direction of the traveling waves Wang et al. (2019a). For simplicity, we set $\Theta=0$ in the following. The details about the derivation of the non-Hermitian Hamiltonian (20) through the Lindblad master equation can be found in appendix B with $\varphi_{a}=0$, despite both $\varphi_{a}$ and $\Gamma$ can have many choices Wang et al. (2019a); Zhang et al. (2017, 2025).
@@
become real-valued, i.e., $E_{\pm}=\pm\sqrt{J^{2}-\gamma_{a}^{2}}$, when $H(t),\mathcal{PT}]=0$ and $J>\gamma_{a}$. Otherwise, they are complex and the Hamiltonian is in the $\mathcal{PT}$-symmetry-broken phase. Whether holding $\mathcal{PT}$-symmetry or not, the system can approach EP under the conditions of $\Delta(t)=0$, $\varphi=\pi/2$, $\gamma_{a}=\gamma_{b}$, and $J=\pm\Gamma$ (when $\varphi_{a}=0$) or $J=\pm\sqrt{\Gamma^{2}+\gamma_{a}^{2}}$ (when $\varphi_{a}=\pi$).
Conventionally, the control protocols proposed for non-Hermitian systems are based on their spectrum characteristics Ashida et al. (2020); Bender and Hook (2024), focusing on $\mathcal{PT}$-symmetry and EPs. In contrast, our UQC theory in Sec. II is essentially proposed in the instantaneous frames that are irrespective to the spectral properties. In the following section, we will demonstrate that an arbitrary state can be perfectly transferred from the cavity mode to magnon mode, regardless of whether the system in the unbroken or broken phases of $\mathcal{PT}$-symmetry and whether EPs are present or not.
@@ III.2 Universal Heisenberg passages of two operators
With Eq. (21), we construct the universal passages for the non-Hermitian cavity-magnonic system, which give rise to nonunitary evolution of the system state and the final state can be automatically normalized. According to Eqs. (42) and (43), the time-dependent ancillary operators of an arbitrary two-mode system can be written as
@@
where the parameters $\theta(t)$ and $\alpha(t)$ manipulate the population and the relative phase of both cavity mode $a$ and magnon mode $b$, respectively. From Eqs. (45) and (46), the unitary transformation $\mathcal{V}(t)$ in Eq. (6) which acts as $\mathcal{V}^{\dagger}(t)\mu_{1}(t)\mathcal{V}(t)\rightarrow\mu_{1}(0)$ and $\mathcal{V}^{\dagger}(t)\mu_{2}(t)\mathcal{V}(t)\rightarrow\mu_{2}(0)$, can take the form of
@@
It is interesting to find that the preceding transformation from the time-dependent ancillary operators to the time-independent ones can also be formally implemented by $V_{\theta}(t)V_{\alpha}(t)]^{\dagger}\mu_{k}(t)V_{\theta}(t)V_{\alpha}(t)\rightarrow\mu_{k}(0)$, $k=1,2$, despite that $V_{\alpha}(t),V_{\theta}(t)]\neq 0$.
@@
In parallel, for the system dynamics governed by the Hermitian conjugate $H^{\dagger}(t)$ in the dual space, the same conditions in Eq. (27) can activate the ancillary operator $\mu_{2}^{\dagger}(t)$ as the nonadiabatic passage. In particular, the time evolution takes the form of Eq. (19), where the global phase $\dot{f}_{2}^{*}(t)=i[\gamma_{a}\exp(-i\varphi_{a})+\gamma_{b}]-\dot{f}_{1}^{*}(t)$ with $\dot{f}_{1}^{*}(t)$ the complex conjugate of $\dot{f}_{1}(t)$ given by Eqs. (29) and (30). Similar to $\mu_{1}^{\dagger}(t)$, a flexible and perfect state transfer can be implemented along the passage $\mu_{2}^{\dagger}(t)$ under appropriate choices of $\theta(t)$ and $\alpha(t)$. The probability conservation can also be ensured in the end of the evolution.
@@ IV.1 Perfect state transfer
We first consider the Hamiltonian in the unbroken phase of $\mathcal{PT}$-symmetry. In this case, $\varphi_{a}=\pi$, $\gamma_{a}=\gamma_{b}$, $\Gamma=0$, and $\Delta(t)=0$. Consequently, the constraint on $J(t)$ in Eq. (27) becomes $J(t)=\dot{\theta}(t)/\sin[\varphi+\alpha(t)]$. Using Eqs. (23), (24), and (28), the system passage can evolve as $\mu_{1}^{\dagger}(0)=a^{\dagger}\rightarrow e^{if_{r}(\tau)}\mu_{1}^{\dagger}(\tau)=e^{if_{r}(\tau)}b^{\dagger}$ under the setting of $\alpha(t)=0$ and
@@
Assume that the cavity mode and magnon mode are initially prepared in the Fock state $|n=5\rangle$ and the vacuum state $|0\rangle$, respectively, i.e.,
$|\psi(0)\rangle=|5\rangle_{a}|0\rangle_{b}=\frac{(a^{\dagger})^{5}}{\sqrt{5!}}|0\rangle_{a}|0\rangle_{b}=\frac{[\mu_{1}^{\dagger}(0)]^{5}}{\sqrt{5!}}|0\rangle_{a}|0\rangle_{b}.$ (32)
It will be finally transformed to be
$\frac{[\mu_{1}^{\dagger}(\tau)]^{5}}{\sqrt{5!}}|0\rangle_{a}|0\rangle_{b}=\frac{(b^{\dagger})^{5}}{\sqrt{5!}}|0\rangle_{a}|0\rangle_{b}=|0\rangle_{a}|5\rangle_{b}=|\psi(\tau)\rangle$ (33)
up to an irrelevant phase, demonstrating a perfect Fock-state transfer from the cavity mode to the magnon mode.
Under the non-unique condition in Eq. (31), we find that the probability conservation of the system state can be restored at the end of the evolution in the presence or in the absence of exceptional points by appropriately choosing the gain or loss rates and the dissipative coupling strength. For the Hamiltonian of $\mathcal{PT}$-symmetry, the imaginary part of the global phase in Eq. (30) is reduced to be
@@
To ensure the vanishing of the time integral over $\dot{f}_{i}(t)$ in Eq. (34), one can choose the gain or loss rate and dissipative coupling strength as
@@
given $\theta(t)$ a linear function of time in Eq. (31). Here the positive factor $\lambda$ scales the magnitudes of these rates or strength.
@@
Figure 2 demonstrates perfect nonadiabatic passage $\mu_{1}^{\dagger}(t)$, i.e., $\mathcal{F}_{5,0}(0)=1$ and $\mathcal{F}_{0,5}(\tau)=1$, under a $\mathcal{PT}$-symmetric Hamiltonian. It is found that for both evolutions of (a) avoiding and (b) crossing EPs, the Fock states with conserved excitations other than the initial and target states are scarcely populated. For example, in Fig. 2(a), we have at most $\mathcal{F}_{4,1}(0.19\tau)=0.073$, $\mathcal{F}_{3,2}(0.40\tau)=0.031$, $\mathcal{F}_{2,3}(0.62\tau)=0.032$, and $\mathcal{F}_{1,4}(0.82\tau)=0.074$. The total fidelity $\sum_{n}\mathcal{F}_{n}\equiv\sum_{n_{1},n_{2}}\mathcal{F}_{n_{1},n_{2}}$, being equivalent to the trace of the two-mode system ${\rm Tr}[|\psi(t)\rangle\langle\psi(t)|]$, can represent the probability conservation or non-conservation under a non-Hermitian Hamiltonian. In both Figs. 2(a) and (b), it is found that $\sum_{n}\mathcal{F}_{n}(0<t<\tau)<1$ due to the gain effect of the cavity mode and the loss effect of the magnon mode. Nevertheless, $\sum_{n}\mathcal{F}_{n}(t=\tau)$ always converges to unit at the end of the passage. In Figs. 2(c) and (d), we present the dynamics of real and imaginary parts of eigenenergies $E_{\pm}$ due to Eq. (22), relevant to the avoiding and crossing EPs in the parametric setting, respectively. The curves in Fig. 2(c) verify that the system dynamics avoids the coalescence of eigenenergies and eigenstates. Figure 2(d) indicates that the system crosses EPs when $t=0.18\tau$ and $t=0.82\tau$. It is hardly to find a clear connection between the perfect state transfer and the existence of the exceptional points of the system spectrum.
@@
Similar to Figs. 2(a) and (b), Figs. 3(a) and (b) demonstrate the fidelity dynamics $\mathcal{F}_{n_{1},n_{2}}$ of various five-excitation product-states and their summation $\sum_{n}\mathcal{F}_{n}=\sum_{n_{1},n_{2}}\mathcal{F}_{n_{1},n_{2}}$ within the broken phase of $\mathcal{PT}$-symmetry in the absence and presence of EPs, respectively. Here we set $\varphi_{a}=0$ and $\gamma_{a}=\gamma_{b}$. Then the imaginary part of the global phase in Eq. (30) becomes
@@
One can check that the gain or loss rate and dissipative coupling strength in Eq. (35) are still applicable to neutralize the imaginary phase under the parametric setting in Eq. (31) for $\theta(t)$.
In both Figs. 3(a) and (b), it is found that the initial population on $|5\rangle_{a}|0\rangle_{0}$ can be completely transferred to $|0\rangle_{a}|5\rangle_{b}$, even when the other five-excitation states could be temporally yet significantly populated during the passage. The Fock states with more excitations of the target (magnon) mode are sequently populated until $\mathcal{F}_{0,5}(\tau)=1$ at the end of the time evolution. For $0<t<\tau$, the summation of individual populations is not conserved, i.e., $\sum_{n}\mathcal{F}_{n}<1$ or $\sum_{n}\mathcal{F}_{n}>1$, due to the loss effects of both the cavity and magnon modes and the dissipative coupling between them. The peak value of the population summation is found to be $\sum_{n}\mathcal{F}_{n}=1.32$ when $t=0.72\tau$ in Fig. 3(a), around which the dissipative coupling between the cavity and magnon modes dominates the loss of both modes. In Fig. 3(a) with no EPs, we have $\mathcal{F}_{4,1}(0.22\tau)=0.35$, $\mathcal{F}_{3,2}(0.34\tau)=0.34$, $\mathcal{F}_{2,3}(0.45\tau)=0.40$, $\mathcal{F}_{1,4}(0.61\tau)=0.55$, and $\mathcal{F}_{0,5}(0.90\tau)=1.07$. And in Fig. 3(b) with EPs, we have $\mathcal{F}_{4,1}(0.35\tau)=0.24$, $\mathcal{F}_{3,2}(0.50\tau)=0.32$, $\mathcal{F}_{2,3}(0.61\tau)=0.46$, $\mathcal{F}_{1,4}(0.72\tau)=0.74$, and $\mathcal{F}_{0,5}(0.89\tau)=1.40$. Under the conditions in Eq. (35), the state norm becomes unit at the end of the passage. Again the associated evolutions of the real and imaginary parts of the eigenenergies are presented in Figs. 3(c) and (d), respectively. As shown in Fig. 3(c), the eigenenergies do not coalesce during the whole passage, confirming the avoidance of EPs. In Fig. 3(d), an EP occurs at $t=0.27\tau$, also being irrelevant to the system dynamics.
@@
Far beyond the Fock state, the universal Heisenberg passage can unravel perfect transfer of diverse states. In Fig. 4(a), the cavity and magnon modes are initially prepared in the binomial code state $|\phi_{b}\rangle=(\sqrt{3}|2\rangle+|6\rangle)/2$ Michael et al. (2016) and the vacuum state $|0\rangle$, respectively. The performance of our protocol can be evaluated by $\mathcal{F}_{\phi_{b},0}=|\langle\phi_{b}|\langle 0|\psi(t)\rangle|^{2}$ and $\mathcal{F}_{0,\phi_{b}}=|\langle 0|\langle\phi_{b}|\psi(t)\rangle|^{2}$. It is found that the binomial code state in the cavity mode can be perfectly transferred to the magnon mode with a unit fidelity. Along the passage, the non-Hermitian system during $0<t<\tau$ is not subject to the probability conservation. For example, we have ${\rm Tr}(\rho)=0.631$ when $t=0.234\tau$ and ${\rm Tr}(\rho)=1.63$ when $t=0.782$. However, the probability conservation is restored as ${\rm Tr}(\rho)=1$ when $t=\tau$. Similarly, in Figs. 4(b), (c), and (d) for the cavity mode initially prepared as the coherent state $|\alpha\rangle$ with $\alpha=5$, the cat state $|{\rm cat}\rangle=(|\alpha\rangle+|-\alpha\rangle)/\sqrt{2}$ with $\alpha=5$, and the thermal state $\rho_{\rm th}$ with $\bar{n}=5$, respectively, it is found that all of them can be perfectly transferred to the target mode. Note for the thermal state, $\mathcal{F}_{{\rm thermal},0}$ and $\mathcal{F}_{0,{\rm thermal}}$ are not complementary to each other in any situation.
In parallel, a quantum state initially prepared in the magnon mode can be perfectly transferred to the cavity mode by activating the passage $\mu_{2}^{\dagger}(t)$ in the dual space. It can be realized under almost the same conditions as for the passage $\mu_{1}^{\dagger}(t)$, except that $\varphi=\pi/2$ is replaced with $\varphi=-\pi/2$.